# The nature of mathematical models

### v1

@@ 2.2 Geometry of Estimation in Case Space ($\mathbb{R}^{n}$)
We summarize here the standard geometrical interpretation of Ordinary Least Squares parameter estimation. While none of the material in this subsection is new (e.g. see Seber [1, 2]), the following description should help to clarify the correspondence of this with subsequently presented material.
When talking about modeling in an applied setting we are typically discussing the “usual representation”
@@
where a dependent variable $x$ is shown to depend upon some independent predictor $u$ (often indicating time in dynamical systems) via some functional form and some $q$-dimensional parameter $\boldsymbol{\theta}$. This is typically graphed showing curves for $x$ as a function of $u$, different curves corresponding to different values of $\boldsymbol{\theta}$. Since by changing the parameter value the prediction $x(u,\boldsymbol{\theta})$ obviously changes, this clearly leads to thinking about estimation as adapting the predicted curve by minimizing some functional of the "errors" (differences between predicted and observed values of $x$), yielding    $\hat{\boldsymbol{\theta}}$    as the $argmin$ of the functional.
However the values $u$ takes are fixed (by design or by sampling), and we may drop $u$ from notation, since what we are really interested in is the relationship between parameter and prediction. Also, we may indicate with $\boldsymbol{x}$ the set of predicted values that $x$ takes for each of $n$ experimental units (corresponding to a given set of predictors) as $\boldsymbol{\theta}$ varies:
@@
and we may also indicate with $\boldsymbol{x}^{o}$ (observed) the actually measured or observed values of $x$ in the $n$ experimental units. We find thus useful to work in ($n$-dimensional) $\boldsymbol{x}$-space or case space, $\mathbb{R}^{n}$. Case space is therefore defined as the $n$-dimensional Euclidean space where each axis represents one of the experimental units, and where along each axis we may identify both the actually observed value for that unit and any predicted value for that unit (depending on the chosen $\boldsymbol{\theta}$ parameter value). In other words, case space is the $n$-dimensional space where all possible values of $\boldsymbol{x}$, either measured or predicted, referring to the $n$ xperimental units can be represented. The appeal of working in case space consists in an immediate visualization of the geometry of the problem: the whole observed sample realization is a single point $\boldsymbol{x}^{o}$ in $n$-dimensional case space, and we may trace a (generally nonlinear) $q$-dimensional prediction surface $\mathcal{S}$ as the set of values that the predicted $\boldsymbol{x}$ can take as $\boldsymbol{\theta}$ varies. Indeed, if $\boldsymbol{\theta}\in\mathbb{R}^{q}$ and $x(\boldsymbol{\theta})$ is sufficiently regular then the prediction surface $\mathcal{S}$ is a $q$-dimensional manifold embedded in case space ($\mathbb{R}^{n}$), having as a global chart map the canonical map $\boldsymbol{x}=\boldsymbol{x}(\boldsymbol{\theta})\in\mathbb{R}^{n}\rightarrow\boldsymbol{\theta}\in\mathbb{R}^{q}$, everywhere.
@@
  • we have the Lebesgue measure $\lambda$ defined on $\mathcal{B}$ with values in +:=$\mathbb{R}^{+}:=\{r\in\mathbb{R},r\geqslant 0\}$, making $(\mathbb{R}^{n},\mathcal{B},\lambda)$ into a measure space
  • we can endow $\mathbb{R}^{n}$ with a norm, the usual euclidean norm $|\boldsymbol{x}|:=\sqrt{\sum_{i=1}^{n}x_{i}^{2}}$
  • with respect to $|\boldsymbol{x}|$ the vector space $\mathbb{R}^{n}$ is complete, hence Banach
  • we can define on $\mathbb{R}^{n}$ an inner product compatible with the norm, $\langle\boldsymbol{x},\boldsymbol{y}\rangle:=\sqrt{\sum_{i=1}^{n}x_{i}y_{i}}$, making $\mathbb{R}^{n}$ Hilbert
@@ 2.2.1 Linear relationship in case space: geometry
  • ${\boldsymbol{\theta}}$ is a $q$-dimensional parameter
  • $\boldsymbol{x}=\boldsymbol{x}(\boldsymbol{\theta})$ is a function mapping (in this case linearly) the set   $\Theta$   of allowable parameter values onto $\mathbb{R}^{n}$, $\boldsymbol{x}:\Theta\rightarrow\mathbb{R}^{n}$
  • $\mathcal{S}$ (the prediction surface) is the set of possible values taken by the function $\boldsymbol{x}(\boldsymbol{\theta})$; in this linear case it is a $q$-dimensional subspace of $\mathbb{R}^{n}$
  • since we assume by hypothesis the relationship given by Eq.3, then $\boldsymbol{x}^{*}$ is the state, the supposedly true value of the observed phenomenon, necessarily lying on the supposedly true expectation surface, at a position indexed by the supposedly true parameter value $\boldsymbol{\theta}^{*}$
@@
  • $\hat{\boldsymbol{x}}$ is our best estimate for the state, the point on $\mathcal{S}$ closest to the observation $\boldsymbol{x}^{o}$, indexed by the parameter value $\hat{\boldsymbol{\theta}}$; notice that it makes sense of speaking about closeness since, as mentioned before, $\mathbb{R}^{n}$ is normed; also, we are arbitrarily taking minimum distance between $\boldsymbol{x}^{o}$ and $\mathcal{S}$ as the basis for our estimation (OLS)
  • $\boldsymbol{{\varepsilon}}=\boldsymbol{x}^{o}-\boldsymbol{x}^{*}$ is the error, the difference between the observation and the true state; since we do not know $\boldsymbol{x}^{*}$ we do not know $\boldsymbol{\varepsilon}$, even though we might make hypotheses about how it should be distributed
  • $\boldsymbol{e}=\boldsymbol{x}^{o}-\hat{\boldsymbol{x}}$ is the residual, the difference between the observation $\boldsymbol{x}^{o}$ and our best guess $\hat{\boldsymbol{x}}$ given the observation and the assumed relationship
  • ${\boldsymbol{\xi}}=\hat{\boldsymbol{x}}-\boldsymbol{x}^{*}$ is the flaw (newly introduced terminology, a name for this object could not be found in the standard literature), the assessment mistake we would make by placing our best guess at $\hat{\boldsymbol{x}}$ instead of at $\boldsymbol{x}^{*}$
@@ 2.2.2 Linear relationship in case space: estimation by projection
Given the above geometry, we can now describe a possible procedure for estimating $\boldsymbol{\theta}$. Notice first that, since $\mathcal{S}$ is a (nonempty) subspace of $\mathbb{R}^{n}$, it is convex and a Hilbert space in its own right. From the Hilbert projection theorem[13] a unique projection operator $\mathcal{P}_{\mathcal{S}}:\mathbb{R}^{n}\rightarrow\mathcal{S}$ can be defined. In fact, if it were $\boldsymbol{x}=\boldsymbol{\theta}U$ for some matrix $U$, we could write $\mathcal{P}_{\mathcal{S}}=U(U^{T}U)^{-1}U^{T}$. Since this projection operator satisfies (always by Hilbert projection theorem)
$\|\boldsymbol{x}-\mathcal{P}_{\mathcal{S}}\boldsymbol{x}\|<\|\boldsymbol{x}-\boldsymbol{y}\|\,,\,\forall\boldsymbol{y}\in\mathcal{S},\boldsymbol{y}\neq\mathcal{P}_{\mathcal{S}}\boldsymbol{x}\,,$
we have that $\hat{\boldsymbol{\theta}}=\underset{\boldsymbol{\theta}\in\Theta}{\arg\min}(||\boldsymbol{x}(\boldsymbol{\theta})-\boldsymbol{x}^{o}||^{2})$ ,  where $\boldsymbol{\Theta}$ is the admissible set of values of $\boldsymbol{\theta}$.
The projection $\hat{\boldsymbol{x}}=\mathcal{P}_{\mathcal{S}}\boldsymbol{x}^{o}$ of the observed sample $\boldsymbol{x}^{o}$ onto the prediction surface $\mathcal{S}$ thus identifies the unique point $\hat{\boldsymbol{x}}\in\mathcal{S}$ which is closest to $\boldsymbol{x}^{o}$ in the OLS sense. The value $\hat{\boldsymbol{\theta}}$ such that $\hat{\boldsymbol{x}}=\boldsymbol{x}(\hat{\boldsymbol{\theta}})$ is the desired (point) parameter estimate.
@@
  • observations are determined by the true relationship, design variable(s) and the supposedly true parameter value $\theta^{*}$, plus error
@@
    asymptotic confidence regions of $\hat{\boldsymbol{x}}=\mathcal{P}_{\mathcal{S}}\boldsymbol{x}^{o}$ on $\mathcal{S}$, hence of $\hat{\boldsymbol{\theta}}$ can be obtained [1].
@@
  • under a rotation of the basis for $\mathbb{R}^{n}$ we can obtain a new basis with the first $q$ elements spanning $\mathcal{S}$ and with the remaining $(n-q)$ elements orthogonal to $\mathcal{S}$;
@@
    $\xi_{j}=0\,\,\forall j\in\{q+1,\dots,n\}$   and   $e_{i}=0\,\,\forall i\in\{1,\dots,q\}$, while
@@
  • $\frac{\|\boldsymbol{\xi}\|^{2}}{\sigma^{2}}=\sum_{j=1}^{n}\,\frac{\xi_{j}^{2}}{\sigma^{2}}=\sum_{j=1}^{q}\,\left(\frac{\xi_{j}}{\sigma}\right)^{2}$,   where   $\frac{\xi_{j}}{\sigma}\sim\mathcal{N}(0,1)$,   so
@@
    $\frac{\|\boldsymbol{e}\|^{2}}{\sigma^{2}}\sim\chi_{n-q}$  (Chi-square with $(n-q)$ degrees of freedom)
  • $\frac{\|\boldsymbol{\xi}\|^{2}}{\sigma^{2}}/\frac{\|\boldsymbol{e}\|^{2}}{\sigma^{2}}=\frac{\|\boldsymbol{\xi}\|^{2}}{\|\boldsymbol{e}\|^{2}}\sim\textit{F}_{q,(n-q)}$  (Fisher with $q$ and $(n-q)$ degrees of freedom)
and we can compute a $(1-\alpha)$ confidence region for $\|\boldsymbol{\xi}\|$ !
In other words: under the stated hypotheses ($\varepsilon_{i}$ i.i.d. $\sim\mathcal{N}(0,\sigma^{2})$), taking into account the dimension $n$ of case space as well as the dimension $q$ of the prediction surface $\mathcal{S}$, by observing the size of the residual from $\mathcal{S}$, $\|\boldsymbol{e}\|$, we have information on the likely size of the flaw along $\mathcal{S}$, $\|\boldsymbol{\xi}\|$ !!!
@@ 2.2.3 Nonlinear relationship in case space
Figure 2: The geometry of a nonlinear relationship in case space
By referring to Fig.2 we see that a construction, similar to what we have provided for the linear case, does not actually hold any more. The fact is that in this context $\mathcal{S}$ is not a subspace of $\mathbb{R}^{n}$ and in general it may not even be convex, hence Hilbert projection theorem does not apply and a unique projection operator $\mathcal{P}_{\mathcal{S}}:\mathbb{R}^{n}\rightarrow\mathcal{S}$ cannot be defined.
We may however look for a value $\hat{\boldsymbol{\theta}}=\underset{\boldsymbol{\theta}\in\boldsymbol{\Theta}}{\arg\min}(||\boldsymbol{x}(\boldsymbol{\theta})-\boldsymbol{x}^{o}||)$ ,  where again $\boldsymbol{\Theta}$ is the admissible set of values of $\boldsymbol{\theta}$, same as for the linear case.
The value $\hat{\boldsymbol{\theta}}$ such that $\hat{\boldsymbol{x}}=\boldsymbol{x}(\hat{\boldsymbol{\theta}})$, the desired (point) parameter estimate, is typically found numerically, by minimization of $\|\boldsymbol{e}(\boldsymbol{\theta})\|$.
We would like to be able to make at least approximate confidence statements about the parameter estimate in the nonlinear case as well. To do so, and in the case that the (local) minimum we have found is isolated, we might wish to use a local linear approximation $T_{\hat{\boldsymbol{x}}}\mathcal{S}$ to the prediction surface around the estimated state $\hat{\boldsymbol{x}}$. Notice that $T_{\hat{\boldsymbol{x}}}\mathcal{S}$ is not in general a subspace, but rather an affine space.
The Hilbert projection theorem however extends naturally to affine subspaces, the key observation being that the problem can be reduced to projecting onto a linear subspace by translating the space appropriately. The projection onto an affine subspace is unique and satisfies an orthogonality condition relative to the associated direction subspace, see Appendix 5.1 for details. In this way, if an isolated minimum of $\|\boldsymbol{e(\boldsymbol{\theta})}\|$ at $\hat{\boldsymbol{x}}$ exists, then we can define a (locally unique) projection operator    $\mathcal{P}_{T_{\hat{\boldsymbol{x}}}\mathcal{S}}:\mathbb{R}^{n}\rightarrow T_{\hat{\boldsymbol{x}}}\mathcal{S}$,    $\mathcal{P}_{T_{\hat{\boldsymbol{x}}}\mathcal{S}}:\boldsymbol{x}^{o}\mapsto\hat{\boldsymbol{x}}$
@@
  • observations are determined by the true relationship, design variable(s) and the supposedly true parameter value $\theta^{*}$, plus error
@@
we have that [2, 3]:
@@
    approximate asymptotic confidence regions of $\hat{\boldsymbol{x}}$ on $T_{\hat{\boldsymbol{x}}}\mathcal{S}$, hence of $\hat{\boldsymbol{\theta}}$ can be obtained: such regions would not be too different from the correct asymptotic confidence regions of $\hat{\boldsymbol{x}}$ on $\mathcal{S}$ if the curvature of $\mathcal{S}$ at $\hat{\boldsymbol{x}}$ is not too large [14].

### v2

@@ 2.2 Geometry of Estimation in Case Space ($\mathbb{R}^{n}$)
We summarize here the standard geometrical interpretation of Ordinary Least
Squares parameter estimation. While none of the material in this section is new[Bates1988, Seber1989, Seber2003], the following description should help clarify the correspondence of this with subsequently presented material.
When talking about a model in an applied setting we typically employ the “usual representation”
@@
where a dependent variable $x$ is assumed to depend upon some independent predictor $u$ (often indicating time in dynamical systems) via some functional form and some $q$-dimensional parameter $\boldsymbol{\theta}$. This relationship is typically graphed showing curves of $x$ as a function of $u$, different curves corresponding to different values of $\boldsymbol{\theta}$. Since by changing the parameter value the prediction $x(u,\boldsymbol{\theta})$ obviously changes, this clearly leads to thinking about estimation as adapting the predicted curve by minimizing some functional of the “errors” (differences between predicted and observed values of $x$), yielding    $\hat{\boldsymbol{\theta}}$    as the  $\operatorname*{arg\,min}$  of the functional.
However, the values $u$ takes are fixed (by design or by sampling), and we may drop $u$ from notation, since what we are really interested in is the relationship between parameter and prediction. Also, we may indicate with $\boldsymbol{x}$ the set of predicted values that $x$ takes for each of $n$ experimental units (corresponding to a given set of predictors) as $\boldsymbol{\theta}$ varies:
@@
[…]
@@
  • we have the Lebesgue measure $\lambda$ defined on $\mathcal{B}$ with values in $\mathbb{R}^{+}:=\{r\in\mathbb{R},r\geqslant 0\}$, making $(\mathbb{R}^{n},\mathcal{B},\lambda)$ into a measure space
  • we can endow $\mathbb{R}^{n}$ with a norm, the usual euclidean norm $\|\boldsymbol{x}\|:=\sqrt{\sum_{i=1}^{n}x_{i}^{2}}$
  • with respect to $\|\boldsymbol{x}\|$ the vector space $\mathbb{R}^{n}$ is complete, hence Banach
  • we can define on $\mathbb{R}^{n}$ an inner product compatible with the norm,  $\langle\boldsymbol{x},\boldsymbol{y}\rangle:=\sum_{i=1}^{n}x_{i}y_{i}$,  making $\mathbb{R}^{n}$ Hilbert
@@ 2.2.1 Linear relationship in case space: geometry
  • $\boldsymbol{\theta}$} is a $q$-dimensional parameter, $\boldsymbol{\theta}\in\Theta\subset\mathbb{R}^{q}$
  • $\boldsymbol{x}=\boldsymbol{x}(\boldsymbol{\theta})$ is a function mapping (in this case linearly) the set   $\Theta$   of allowable parameter values into $\mathbb{R}^{n}$, $\boldsymbol{x}:\Theta\rightarrow\mathbb{R}^{n}$
  • $\mathcal{M}_{\boldsymbol{x}}$} (the prediction surface) is the set of possible values taken by the function $\boldsymbol{x}(\boldsymbol{\theta})$; in this linear case it is a $q$-dimensional subspace of $\mathbb{R}^{n}$
  • since we assume by hypothesis the relationship given by Eq.3, then $\boldsymbol{x}^{*}$ is the state, the supposedly true value of the observed phenomenon, necessarily lying on the supposedly true prediction surface, at a position indexed by the supposedly true parameter value $\boldsymbol{\theta}^{*}$
@@
  • $\hat{\boldsymbol{x}}$ is our best estimate for the state, the point on $\mathcal{M}_{\boldsymbol{x}}$} closest to the observation $\boldsymbol{x}^{o}$, indexed by the parameter value $\hat{\boldsymbol{\theta}}$; notice that it makes sense to speak about closeness since, as mentioned before, $\mathbb{R}^{n}$ is normed; also, we are arbitrarily taking minimum distance between $\boldsymbol{x}^{o}$ and $\mathcal{M}_{\boldsymbol{x}}$} as the basis for our estimation (OLS)
  • $\boldsymbol{{\varepsilon}}^{o}=\boldsymbol{x}^{o}-\boldsymbol{x}^{*}$ is the error made by observing $\boldsymbol{x}^{o}$ instead of $\boldsymbol{x}^{*}$, the difference between the observation and the true state; it is the actual realization of the random variable $\boldsymbol{\varepsilon}$}: since we do not know $\boldsymbol{x}^{*}$ we do not know ${\boldsymbol{\varepsilon}}^{o}$, even though we might make hypotheses about how $\boldsymbol{\varepsilon}$} should be distributed
  • $^{\boldsymbol{e}}^{o}=\boldsymbol{x}^{o}-\hat{\boldsymbol{x}}$ is the residual, the observed difference between the observation $\boldsymbol{x}^{o}$ and our best guess $\hat{\boldsymbol{x}}$; once again, ${\boldsymbol{e}}^{o}$ is the actual realization of the random variable $\boldsymbol{e}$}
  • ${\boldsymbol{\xi}}^{o}=\hat{\boldsymbol{x}}-\boldsymbol{x}^{*}$ is the actual flaw (we cannot qualify it as “observed” flaw, since we do not really observe it)  [^1] The term “flaw” is here introduced, since a name for $\boldsymbol{\xi}$ could not be found in the standard literature, the assessment mistake we would make by placing our best guess at $\hat{\boldsymbol{x}}$ instead of at $\boldsymbol{x}^{*}$ and, as before, it is the realization of the random variable $\boldsymbol{\xi}$.
@@ 2.2.2 Linear relationship in case space: estimation by projection
Given the above geometry, we can now describe a possible procedure for estimating $\boldsymbol{\theta}$. Notice first that, since $\mathcal{M}_{\boldsymbol{x}}$} is a (nonempty) subspace of $\mathbb{R}^{n}$, it is convex and a Hilbert space in its own right. From the Hilbert projection theorem[Rudin1987] a unique projection operator $\mathcal{P}_{\mathcal{M}_{\boldsymbol{x}}}:\mathbb{R}^{n}\rightarrow{\mathcal{M}_{\boldsymbol{x}}}$ can be defined. In fact, if it were $\boldsymbol{x}=U\boldsymbol{\theta}$ for some matrix $U$, we could write $\mathcal{P}_{\mathcal{M}_{\boldsymbol{x}}}=U(U^{\intercal}U)^{-1}U^{\intercal}$. Since this projection operator satisfies (always by the Hilbert projection theorem)
$\|\boldsymbol{x}-\mathcal{P}_{\mathcal{M}_{\boldsymbol{x}}}\boldsymbol{x}\|<\|\boldsymbol{x}-\boldsymbol{y}\|\,,\,\forall\boldsymbol{y}\in{\mathcal{M}_{\boldsymbol{x}}},\boldsymbol{y}\neq\mathcal{P}_{\mathcal{M}_{\boldsymbol{x}}}\boldsymbol{x}\,$,
[…]
@@
  • observations are determined by the true relationship, design variable(s) and the supposedly true parameter value $\boldsymbol{\theta}^{*}$, plus error
@@
    asymptotic confidence regions of $\hat{\boldsymbol{x}}=\mathcal{P}_{\mathcal{M}_{\boldsymbol{x}}}\boldsymbol{x}^{o}$ on $\mathcal{M}_{\boldsymbol{x}}$}, hence of $\hat{\boldsymbol{\theta}}$ can be obtained [Seber2003].
@@
  • under a rotation of the basis for $\mathbb{R}^{n}$ we can obtain a new basis with the first $q$ elements spanning $\mathcal{M}_{\boldsymbol{x}}$} and with the remaining (nq)(n-q) elements orthogonal to $\mathcal{M}_{\boldsymbol{x}}$};
@@
    $e_{j}=0\,\,\forall j\in\{1,\dots,q\}$   and   $\xi_{i}=0\,\,\forall i\in\{q+1,\dots,n\}$ , while
@@
  • $\frac{\|\boldsymbol{\xi}\|^{2}}{\sigma^{2}}=\sum_{i=1}^{n}\,\frac{\xi_{i}^{2}}{\sigma^{2}}=\sum_{j=1}^{q}\,\left(\frac{\xi_{j}}{\sigma}\right)^{2}$,   where   $\frac{\xi_{j}}{\sigma}\sim\mathcal{N}(0,1)$,   so
@@
    $\frac{\|\boldsymbol{e}\|^{2}}{\sigma^{2}}\sim\chi_{n-q}$  (Chi-square with (nq)(n-q) degrees of freedom)
  • $\frac{\|\boldsymbol{\xi}\|^{2}}{\sigma^{2}}/\frac{\|\boldsymbol{e}\|^{2}}{\sigma^{2}}=\frac{\|\boldsymbol{\xi}\|^{2}}{\|\boldsymbol{e}\|^{2}}\sim\textit{F}_{q,(n-q)}$  (Fisher with $q$ and (nq)(n-q) degrees of freedom)
and we can compute a $(1-\alpha)$ confidence region for $\|\boldsymbol{\xi}\|$.
In other words: under the stated hypotheses ($\varepsilon_{i}$ i.i.d. $\sim\mathcal{N}(0,\sigma^{2})$), taking into account the dimension $n$ of case space as well as the dimension $q$ of the prediction surface $\mathcal{M}_{\boldsymbol{x}}$}, by observing the size of the residual from $\mathcal{M}_{\boldsymbol{x}}$}, $\|\boldsymbol{e}\|$, we have information on the likely size of the flaw along $\mathcal{M}_{\boldsymbol{x}}$}, $\|\boldsymbol{\xi}\|$.
@@ 2.2.3 Nonlinear relationship in case space
Figure 2: Nonlinear relationship in case space: original geometry
By referring now to Fig.2 at page 2 we see that a construction, similar to what was provided for the linear case, does not allow us to draw the same conclusions. The fact is that in this context $\mathcal{M}_{\boldsymbol{x}}$} is not a subspace of $\mathbb{R}^{n}$ and in general it may not even be convex, hence the Hilbert projection theorem does not apply and a unique projection operator $\mathcal{P}_{\mathcal{M}_{\boldsymbol{x}}}:\mathbb{R}^{n}\rightarrow{\mathcal{M}_{\boldsymbol{x}}}$ cannot be defined.
[…]
We would like to be able to make at least approximate confidence statements about the parameter estimate in the nonlinear case as well. To do so, and in case an isolated local minimum has been found, we might wish to use a local linear approximation $\mathcal{T}_{\hat{\boldsymbol{x}}}={\hat{\boldsymbol{x}}}+T_{\hat{\boldsymbol{x}}}{\mathcal{M}_{\boldsymbol{x}}}$ to the prediction surface $\mathcal{M}_{\boldsymbol{x}}$} around the estimated state $\hat{\boldsymbol{x}}$. Notice that  $\mathcal{T}_{\hat{\boldsymbol{x}}}$  is not in general a subspace, but rather an affine space.
The Hilbert projection theorem however extends naturally to affine subspaces, the key observation being that the problem can be reduced to projecting onto a linear subspace by translating the space appropriately. The projection onto an affine subspace is unique and satisfies an orthogonality condition relative to the associated direction subspace   $T_{\hat{\boldsymbol{x}}}{\mathcal{M}_{\boldsymbol{x}}}$  , see Appendix 6.1 for details. In this way, if an isolated minimum of $\boldsymbol{x}(\boldsymbol{\theta})-\boldsymbol{x}^{o}||^{2}$ at $\hat{\boldsymbol{x}}$ exists, then we can define a (locally unique) projection operator    $\mathcal{P}_{\mathcal{T}_{\hat{\boldsymbol{x}}}}:\mathbb{R}^{n}\rightarrow\mathcal{T}_{\hat{\boldsymbol{x}}}$,    $\mathcal{P}_{\mathcal{T}_{\hat{\boldsymbol{x}}}}:\boldsymbol{x}^{o}\mapsto\hat{\boldsymbol{x}}$ .
@@
  • observations are determined by the true relationship, design variable(s) and the supposedly true parameter value $\boldsymbol{\theta}^{*}$, plus error
@@
we have that [Seber1989, Bates1988]:
@@
    approximate asymptotic confidence regions of $\hat{\boldsymbol{x}}$ on $\mathcal{T}_{\hat{\boldsymbol{x}}}$, hence of $\hat{\boldsymbol{\theta}}$ can be obtained: such regions would not be too different from the correct asymptotic confidence regions of $\hat{\boldsymbol{x}}$ on $\mathcal{M}_{\boldsymbol{x}}$} if the curvature of $\mathcal{M}_{\boldsymbol{x}}$} at $\hat{\boldsymbol{x}}$ is not too large [Panunzi2005].